OpenStudy (anonymous):

Write the equation 2x - 3y = 6 in slope-intercept form.

OpenStudy (anonymous):

thats just rearranging for y

OpenStudy (anonymous):

y=2/3x-2

OpenStudy (anonymous):

so 2x-3y=6 2x=6+3y 2x-6=3y y=(2x-6)/3

OpenStudy (anonymous):

Thanks To Both OF YOu!